OpenStudy (anonymous):

Write a proportion to solve the problem. Then solve the proportion. 2. Seven is 20% of what number?

OpenStudy (anonymous):

7 / .20 = 35

OpenStudy (anonymous):

7 divide by 1/5 = 7*5 = 35